OpenStudy (jozelynw):

Simplify the expression √32+√8−√2 1.)√4•4•2+√4•2−√2 2.)2•2√2+2√2−√2 For the second part where are these numbers coming from. This is from my math class,so the info is correct. I just don't understand how the teacher got these numbers. It just has the numbers not any explanation.

OpenStudy (johnweldon1993):

Well we start with \[\large \sqrt{32} + \sqrt{8} - \sqrt{2}\] think, how else can you write 32? the teacher has pointed out that you can write 32 as the product of 4 times 4 times 2 and the square root of 8....you can write 8 as the product of 4 times 2 so we would have \[\large \sqrt{4\times 4\times 2} + \sqrt{4\times 2} - \sqrt{2}\]

OpenStudy (jozelynw):

Right

OpenStudy (johnweldon1993):

now what has THAT accomplished? all we did was write more numbers but what is different about some of these numbers? Hint...what is the square root of 4?

OpenStudy (jozelynw):

2

OpenStudy (johnweldon1993):

right...so we can simplify Focus on the first square root part here \[\large \sqrt{4\times 4\times 2}\] we can also write this *since it is multiplication inside* as \[\large \sqrt{4} \times \sqrt{4} \times \sqrt{2}\] right? so since we know the square root of = 2 we have \[\large 2 \times 2 \times \sqrt{2}\] or \[\large 4\sqrt{2}\] right?

OpenStudy (jozelynw):

Yes

OpenStudy (johnweldon1993):

And using that same logic we know \[\large \sqrt{4 \times 2} \rightarrow \sqrt{4}\times \sqrt{2} \rightarrow 2\sqrt{2}\]

OpenStudy (jozelynw):

Thank you @johnweldon1993

OpenStudy (johnweldon1993):

No problem!